# Simplifying Subgraph Representation Learning for Scalable Link Prediction

### v3

@@ 1. Introduction
The core issue hampering the deployment of GNNs and SGRLs is their high computational demands on large-scale graphs (e.g., OGB datasets (Hu et al., 2020, 2021)). Many approaches have been proposed for GNNs to relieve this bottleneck, ranging from sampling (Chen et al., 2018; Zeng et al., 2020) to simplification (Wu et al., 2019) techniques. However, these techniques fail to be applied directly in SGRL models, where the issue is even more adverse due to subgraph extractions pertaining to each link. Although recent work has focused on tackling SGRL scalability by sampling (Yin et al., 2022; Louis et al., 2022) or sketching (Chamberlain et al., 2023) approaches, little attention is given to simplification techniques. We focus on improving the scalability in SGRLs by simplifying the underlying training mechanism.
@@
  • We proposed our S3GRL framework to extend simplification methods introduced in graph neural networks to the context of subgraph representation learning.
@@
  • We empirically study and evaluate the performance of our S3GRL instances on 14 small-to-large graph datasets under various evaluation criteria. Our empirical studies and analyses include deep comparisons of our methods against 16 competitive baselines widely utilized in link prediction studies.
  • Our empirical results confirm the efficacy improvements of our S3GRL methods over others while offering multi-fold speed-up in training and inference.
  • Our results also demonstrate a multi-fold reduction in storage requirements: The S3GRL preprocessed datasets are multiple magnitudes smaller than those of other SGRLs (e.g., SEAL (Zhang and Chen, 2018)).
  • We share our findings and framework as an open-source framework [^1] https://github.com/venomouscyanide/S3GRL_OGB, enabling other researchers to devise and test their own S3GRL instances, thereby advancing link prediction research.
@@ 2. Related Work
Recently, a new research direction has emerged in GNNs on scalable training and inference techniques, which mainly focuses on node classification tasks. A common practice is to use different sampling techniques—such as node sampling (Hamilton et al., 2017), layer-wise sampling (Zou et al., 2019), and graph sampling (Zeng et al., 2020)—to subsample the training data to reduce computation. Other approaches utilize simplification techniques, such as removing intermediate non-linearities (Wu et al., 2019; Frasca et al., 2020), to speed up the training and inference. These approaches have been shown to fasten learning of MPGNNs (at the global level) for node and graph classification tasks, and even for link prediction (Pho and Mantzaris, 2022); however, they are not directly applicable to SGRLs which exhibit superior performance for link prediction. In this work, we take the scalability-by-simplification approach for SGRLs (rather than for MPGNNs) by introducing Scalable Simplified SGRL (S3GRL). Allowing diverse definitions of subgraphs and subgraph-level diffusion operators for SGRLs, S3GRL emulates SGRLs in generalization while speeding them up.
@@ 3. Preliminaries
where normalized adjacency matrix $\hat{\mathbf{A}}=\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}$, $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$, $\tilde{\mathbf{D}}$ is the diagonal matrix with $\tilde{D}_{ii}=\sum_{j}\tilde{A}_{ij}$, and $\sigma$ is a non-linearity function (e.g., ReLU). Here, $\mathbf{H}^{(0)}=\mathbf{X}$ and $\mathbf{W}^{(l)}$ is the $l^{th}$ layer’s learnable weight matrix. After $L$ stacked layers of Eq. 1, GCN outputs nodal-representation matrix $\mathbf{Z}=\mathbf{H}^{(L)}$. For any target nodes $T=\{u,v\}$, one can compute its joint representation:
@@
where $\mathbf{z}_{u}$, $\mathbf{z}_{v}$ are $u$’s and $v$’s learned representations in $\mathbf{Z}$, and $\operatorname{pool}$ is a pooling/readout operation (e.g., Hadamard product or concatenation) to aggregate the pairs’ representations. Then, the link probability $p(u,v)$ for the target is given by $p(u,v)=\Omega(\mathbf{q}_{uv})$, where $\Omega$ is a learnable non-linear function (e.g., MLP) transforming the target embedding $\mathbf{q}_{uv}$ into a link probability.
@@ 5.3. Graph Pooling
Non-attr. NS
1,461
2,742
3.75
NA
Power
4,941
6,594
2.67
NA
Yeast
2,375
11,693
9.85
NA
PB
1,222
16,714
27.36
NA
Attributed Cora
2,708
4,488
3.31
1,433
CiteSeer
3,327
3,870
2.33
3,703
PubMed
19,717
37,676
3.82
500
Texas
183
143
1.56
1,703
Wisconsin
251
197
1.57
1,703
OGB Collab
235,868
1,285,465
8.2
128
DDI
4,267
1,334,889
500.5
NA
Vessel
3,538,495
5,345,897
2.4
3
PPA
576,289
30,326,273
73.7
58
Citation2
2,927,963
30,561,187
20.7
128
@@ 5.4. Time Complexities and Storage Requirement
Time Complexity Discussion (PoS vs SEAL). The preprocessing complexity of PoS $O(rnm+rmd)$ is noticeably greater than that of SEAL for both dynamic $O(1)$ and static $O(m)$ modes . However, this higher complexity is manageable in practice by resorting to sampling such as PoS + ScaLed $O(rk^{2}l^{2}+kld)$ or by leveraging distributed computing infrastructures such as Apache Spark. Even without such mitigation strategies, we argue that incurring this relatively high pre-processing cost to gain faster inference and learning offers efficiency in practice. This contrasts SEAL, suffering from a considerably high inference time complexity of $O(Lmdd^{\prime}+d^{\prime 2})$, which depends on graph size (e.g., the number of edges $m$). It is important to note that the computational cost of inference is repetitive and accumulated over multiple training iterations. Our experiments below validate this argument.
@@ 6. Experiments
Baselines. For attributed and non-attributed datasets, we compare our S3GRL models (PoS, $\text{PoS}^{\text{+}}$, and SoP) with 15 baselines belonging to five different categories of link prediction models. Our heuristic benchmarks include common neighbors (CN), Adamic Adar (AA) (Adamic and Adar, 2003), and personalized pagerank (PPR). For message-passing GNNs (MPGNNs), we select GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017) and GAT (Vaswani et al., 2017). Our latent factor (LF) benchmarks are node2vec (Grover and Leskovec, 2016) and Matrix Factorization (Koren et al., 2009). The autoencoder (AE) methods include GAE & VGAE (Kipf and Welling, 2016), adversarially regularized variational graph autoencoder (ARVGA) (Pan et al., 2018) and Graph InfoClust (GIC) (Mavromatis and Karypis, 2021). Our examined SGRLs include SEAL (Zhang and Chen, 2018), GCN+DE (Li et al., 2020), and WalkPool (Pan et al., 2022). For OGB datasets, we compare against BUDDY (Chamberlain et al., 2023) and its baselines (i.e., a subset of our baselines listed above).
@@
Dataset compression of S3GRL vs SEAL. The preprocessing in SGRLs imposes additional disk space requirements to store preprocessed datasets including enclosing subgraphs for each target pair, and their augmented nodal features. Following up our theoretical analyses in Section 1, we intend to study how effective our S3GRL methods are in compressing the preprocessed datasets. Table 7 reports the disk space requirements of S3GRL models vs SEAL and their effectiveness in disk space reduction. The S3GRL has significantly reduced the space requirement, where the maximum is observed in PB in which PoS reduces the 28GB SEAL (preprocessed) training dataset to only 32MB dataset with the 99.88% reduction. For another example, PoS achieves a 99% reduction in PubMed over SEAL: The SEAL training dataset requires 280GB of disk space in comparison to only about 2GB requirement in all S3GRL models.

### v4

@@ 1. Introduction
The core issue hampering the deployment of GNNs and SGRLs is their high computational demands on large-scale graphs (e.g., OGB datasets (Hu et al., 2020, 2021)). Many approaches have been proposed for GNNs to relieve this bottleneck, ranging from sampling (Chen et al., 2018; Zeng et al., 2020) to simplification (Wu et al., 2019) techniques. However, these techniques cannot be directly applied in SGRL models, where the issue is even more adverse due to subgraph extractions pertaining to each link. Although recent work has focused on tackling SGRL scalability by sampling (Yin et al., 2022; Louis et al., 2022) or sketching (Chamberlain et al., 2023) approaches, there is a lack of simplification techniques specifically designed for SGRLs. We focus on improving the scalability in SGRLs by simplifying the underlying training mechanism.
The motivation for our work stems from the increasing need for scalable and efficient link prediction methods in large-scale graphs. Traditional link prediction approaches often struggle with scalability and computational efficiency when applied to these massive datasets. Our Simplified subgraph representation learning approach offers a promising solution by reducing computational complexity while maintaining predictive performance.
@@
  • We introduce the S3GRL framework, extending simplification methods from GNNs to the context of subgraph representation learning, addressing the scalability issue in SGRLs for link prediction.
@@
  • We empirically study and evaluate the performance of our S3GRL instances on 14 small-to-large graph datasets under various evaluation criteria. Our empirical studies and analyses include deep comparisons of our methods against 16 competitive baselines widely utilized in link prediction studies. Our empirical results confirm the efficacy improvements of our S3GRL methods over others while offering multi-fold speed-up in training and inference. Our results also demonstrate a multi-fold reduction in storage requirements: The S3GRL preprocessed datasets are multiple magnitudes smaller than those of other SGRLs (e.g., SEAL (Zhang and Chen, 2018)).
  • We contribute to the research community by sharing our findings and framework as an open-source framework [^1] https://github.com/venomouscyanide/S3GRL_OGB. This enables other researchers to devise and test their own S3GRL instances, fostering further advancements in link prediction research.
@@ 2. Related Work
Recently, a new research direction has emerged in GNNs on scalable training and inference techniques, which mainly focuses on node classification tasks. A common practice is to use different sampling techniques—such as node sampling (Hamilton et al., 2017), layer-wise sampling (Zou et al., 2019), and graph sampling (Zeng et al., 2020)—to subsample the training data to reduce computation. Other approaches utilize simplification techniques (e.g., removing intermediate non-linearities (Wu et al., 2019; Frasca et al., 2020; Saber and Salehi-Abari, 2024)), hashing techniques (Wu et al., 2021), or approximate search techniques (Wang et al., 2022) to speed up the training and inference. These approaches have been shown to fasten learning of MPGNNs (at the global level) for node and graph classification tasks, and even for link prediction (Pho and Mantzaris, 2022; Wang et al., 2022; Wu et al., 2021); however, they are not directly applicable to SGRLs which exhibit superior performance for link prediction. In this work, we take the scalability-by-simplification approach for SGRLs (rather than for MPGNNs) by introducing Scalable Simplified SGRL (S3GRL). Allowing diverse definitions of subgraphs and subgraph-level diffusion operators for SGRLs, S3GRL emulates SGRLs in generalization while speeding them up.
@@ 3. Preliminaries
where normalized adjacency matrix $\hat{\mathbf{A}}=\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}$, $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$, $\mathbf{I}$ is an identity matrix, $\tilde{\mathbf{D}}$ is the diagonal matrix with $\tilde{D}_{ii}=\sum_{j}\tilde{A}_{ij}$, and $\sigma$ is a non-linearity function (e.g., ReLU). Here, $\mathbf{H}^{(0)}=\mathbf{X}$ and $\mathbf{W}^{(l)}$ is the $l^{th}$ layer’s learnable weight matrix. After $L$ stacked layers of Eq. 1, GCN outputs nodal-representation matrix $\mathbf{Z}=\mathbf{H}^{(L)}$. For any target nodes $T=\{u,v\}$, one can compute its joint representation:
@@
where $\mathbf{z}_{u}$, $\mathbf{z}_{v}$ are $u$’s and $v$’s learned representations in $\mathbf{Z}$, and $\operatorname{pool}$ is a pooling/readout operation (e.g., Hadamard product or concatenation) to aggregate the pairs’ representations. Then, the link probability $P_{uv}$ for the target is given by $P_{uv}=\Omega(\mathbf{q}_{uv})$, where $\Omega$ is a learnable non-linear function (e.g., MLP) transforming the target embedding $\mathbf{q}_{uv}$ into a link probability.
@@ 5.3. Graph Pooling
Non-attr. NS 1,461 2,742 3.75 NA
Power 4,941 6,594 2.67 NA
Yeast 2,375 11,693 9.85 NA
PB 1,222 16,714 27.36 NA
Attributed Cora 2,708 4,488 3.31 1,433
CiteSeer 3,327 3,870 2.33 3,703
PubMed 19,717 37,676 3.82 500
Texas 183 143 1.56 1,703
Wisconsin 251 197 1.57 1,703
OGB Collab 235,868 1,285,465 8.2 128
DDI 4,267 1,334,889 500.5 NA
Vessel 3,538,495 5,345,897 2.4 3
PPA 576,289 30,326,273 73.7 58
Citation2 2,927,963 30,561,187 20.7 128
@@ 5.4. Time Complexities and Storage Requirement
Time Complexity Discussion (PoS vs SEAL). The preprocessing complexity of PoS $O(rnm+rmd)$ is noticeably greater than that of SEAL for both dynamic $O(1)$ and static $O(m)$ modes . However, this higher complexity is manageable in practice by resorting to sampling such as PoS + ScaLed $O(rk^{2}l^{2}+kld)$ or by leveraging distributed computing infrastructures such as Apache Spark. Even without such mitigation strategies, we argue that incurring this relatively high pre-processing cost to gain faster inference and training offers efficiency in practice. This contrasts SEAL, suffering from a considerably high inference time complexity of $O(Lmdd^{\prime}+d^{\prime 2})$, which depends on graph size (e.g., the number of edges $m$). It is important to note that the computational cost of inference is repetitive and accumulated over multiple training iterations. Our experiments below validate this argument.
However, for single-instance inference (e.g., test data), S3GRL’s scalability benefits may be less pronounced than SEAL, as both preprocessing and inference are required. In this scenario, the total complexity for PoS becomes $O(rnm+rmd+rdd^{\prime}+d^{\prime 2})$, while for SEAL it is $O(Lmdd^{\prime}+d^{\prime 2})$. Depending on the graph size ($n$ nodes and $m$ edges) and the feature dimensionalities ($d$ and $d^{\prime}$), either method may be faster. Nonetheless, this does not undermine the scalability advantages of S3GRL over SEAL during training.
@@ 6. Experiments
Baselines. To thoroughly evaluate the performance of our S3GRL models (PoS, $\text{PoS}^{\text{+}}$, and SoP), we compare them with a diverse set of 18 baseline methods spanning five categories of link prediction models. The selection of baselines is motivated by the need to cover various techniques, including classic and recent state-of-the-art approaches, to compare comprehensively. Our heuristic benchmarks include common neighbors (CN), Adamic Adar (AA) (Adamic and Adar, 2003), and Personalized PageRank (PPR), which are traditional yet effective and scalable methods for link prediction. For message-passing GNNs (MPGNNs), we select widely used architectures such as GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), and GIN (Xu et al., 2019). These models represent foundational MPGNN techniques commonly used as benchmarks in the literature. Our latent factor (LF) benchmarks are node2vec (Grover and Leskovec, 2016) and Matrix Factorization (MF) (Koren et al., 2009), which are representative of embedding-based methods for link prediction. The autoencoder (AE) methods include GAE and VGAE (Kipf and Welling, 2016), adversarially regularized variational graph autoencoder (ARVGA) (Pan et al., 2018), and Graph InfoClust (GIC) (Mavromatis and Karypis, 2021). These methods are selected for their ability to learn effective representations through unsupervised learning. To ensure inclusion of recent advancements in subgraph representation learning (SGRL), our examined SGRLs include SEAL (Zhang and Chen, 2018), GCN+DE (Li et al., 2020), WalkPool (Pan et al., 2022), and BUDDY (Chamberlain et al., 2023). SEAL is a seminal work in SGRL for link prediction, while GCN+DE incorporates distance encoding for enhanced performance. WalkPool and BUDDY are more recent methods that have shown strong performance on large-scale graphs. For OGB datasets, we compare against BUDDY (Chamberlain et al., 2023) and its baselines (i.e., a subset of our baselines listed above). By selecting these baselines, we aim to provide a comprehensive evaluation of our models against both traditional and cutting-edge methods, addressing the need for up-to-date comparisons.
@@
Dataset compression of S3GRL vs SEAL. The preprocessing in SGRLs imposes additional disk space requirements to store preprocessed datasets including enclosing subgraphs for each target pair, and their augmented nodal features. Our study focuses on the disk space requirements of the final preprocessed datasets, as these represent the long-term storage needs for deploying SGRL models. Following up our theoretical analyses in Section 1, we intend to study how effective our S3GRL methods are in compressing the preprocessed datasets. Table 7 reports the disk space requirements of S3GRL models vs SEAL and their effectiveness in disk space reduction. The S3GRL has significantly reduced the space requirement, where the maximum is observed in PB in which PoS reduces the 28GB SEAL (preprocessed) training dataset to only 32MB dataset with the 99.88% reduction. For another example, PoS achieves a 99% reduction in PubMed over SEAL: The SEAL training dataset requires 280GB of disk space in comparison to only about 2GB requirement in all S3GRL models. [^10] Although we have not included an analysis of the intermediate storage requirements during preprocessing, we believe that the primary concern for practical applications is the disk space required to store the preprocessed datasets, as this impacts the feasibility of deploying and sharing models, especially on systems with limited storage capacity. Furthermore, since our methods process subgraphs sequentially without needing to hold large amounts of data in memory or on disk at once, the intermediate storage requirements are manageable and do not significantly affect the overall resource usage.